# 2Business Informatics Group, TU Wien, Austria

### 19 Dec 2023 v2

@@ 5.1 Sustainability maturity levels
typically seeks a balance between the quality, cost, and delivery time of engineered systems. At this level of maturity, efforts primarily aim at a better understanding of these three factors in either of the two dimensions of the framework in Figure 2, but not in combination of the two. This typically leads to advanced process methods in the end-to-end dimension of the framework in Figure 2. Substantial work has been dedicated to engineering process modeling Pahl and Beitz (1996), assessment Rout et al. (2007), optimization David et al. (2016), and adaptation Seiger et al. (2019). Occasionally, organizations might realize that quality, cost, and time-to-market can be understood through the lens of sustainability, leading to advancement in the sustainability dimension of the framework. For example, sustainability can be interpreted as a property of systems quality Lago et al. (2015), which allows for a sustainability-driven assessment of quality and cost factors. Ensuring technical sustainability Razavian et al. (2014) also allows for reduced costs by proper evolution after the system has been put in place. Currently, sustainability efforts are mostly focused on energy consumption of software-intensive systems Ournani et al. (2020); Palomba et al. (2019) and advanced digital facilities Bellis and Denil (2022), and reducing waste of engineering processes Andersson et al. (2011).
@@
combines elements of advanced process methods and extended sustainability scopes. Most importantly, this level of sustainability maturity allows reasoning about system-level sustainability along the overall engineering process and across process networks. This is demonstrated by sustainability-focused end-to-end methods, such as process scheduling for waste management (Le Hesran et al., 2019). Another example is manufacturing rationalization Mouzon et al. (2007); Gao et al. (2018) enabled by an improved understanding of the end-to-end process network and the resulting improved schedulability. Unfortunately, the stratified nature of sustainability McGuire et al. (2023), hinders the development of methods that would combine process-based longitudinal reasoning with reasoning about sustainability, despite the well-understood benefits (Daoutidis et al., 2016; Barisic et al., 2023). Thus, sustainable systems engineering is an ambitious goal for the majority of organizations currently.
@@
improves on the final limitation of sustainable systems engineering and allows for combining end-to-end concerns with bipartite sustainability. In this combination of concerns, value retention loops are facilitated through the thorough understanding of $(i)$ the network of systems engineering processes, $(ii)$ the sustainability properties of the engineered system, and $(iii)$ the sustainability properties of the employed engineering methods. Circular systems engineering enables assessing and optimizing for systems properties, including sustainability properties across multiple engineering endeavors and allows for finding sustainability trade-offs across different engineered systems. In the most characteristic case, trade-offs can be identified between a system that is upcycled Wegener (2016) into another system—and that, at the active design time or lifetime of the original system. Design for reuse Richardson (2011) and sustainable remanufacturing Jensen et al. (2019) are examples of initiatives that can benefit from a circular systems engineering mindset. Recently, digital twins opened up opportunities for efficient data harvesting and rich decision support based on modeling and simulation Heithoff et al. (2023) and might serve as the enablers of this level of maturity.
@@ 5.2 Sustainability maturation strategies
Maturation strategies from traditional to circular systems engineering can generally follow three pathways as follows.
@@
aims at a better understanding of the value chain and prioritizes improvements in the end-to-end dimension before dedicating efforts to sustainability.
Dahlgaard and Dahlgaard (1999) define operational excellence (OpEx) by the four ‘P’s of excellence: people, partnerships, processes, and products. Covering the end-to-end value chain is of particular importance in achieving operational excellence. Along an operational excellence-driven strategy, a company would develop its internal business capabilities to the level where internal organizations and departments are well-connected through the enterprise process network layer, allowing for efficient flow of information and advanced analytics techniques such as real-time and predictive analytics Oberdorf et al. (2023) that help achieve excellence.
This strategy might be dominant in sectors where sustainability concerns of engineered systems and more importantly, engineering methods are dominated by other functional and extra-functional criteria. Such a maturation strategy can be observed in software engineering where lifecycle models have been researched for decades, but we are just at the beginning of understanding the environmental impacts of computationally demanding methods.
Eventually, an OpEx-driven strategy would implement more and more elaborate sustainability concerns in order to arrive at circular systems engineering.
@@
This strategy might be dominant in sectors where sustainability concerns are of particular importance either due to legislation or social expectations. Pertinent examples include industries with increased environmental impact, such as machine production, cyber-physical systems, or the automotive industry.
@@
aims to balance between the previous two strategies, gradually expanding in end-to-end and sustainability concerns.
@@
This strategy might be the most typical and natural as it allows for quick wins at a particular time and in context, and the refined scopes of iteration minimize strategic risks.
@@ 6 Challenges and Research Opportunities
System complexity often renders the assessment of sustainability and finding trade-offs among functional and sustainability properties a wicked problem Rittel (1967). Such wicked problems are often impossible to solve due to the complex and interconnected nature of the components of systems and the ensuing sub-problems. For example, of the seventeen SDGs formulated by the UN, fourteen directly incorporate socio-economic elements (goals 1-12, 16, and 17), rendering their modeling and analysis a particularly challenging problem. In such a convoluted problem space, modeling, and especially model-based and model-driven engineering techniques Schmidt (2006) are powerful tools in taming complexity Bork et al. (2023). Through the mechanism of abstraction, model-driven techniques foster an understanding of sustainability properties that are known to be stratified (have different meanings at different levels of abstraction) and multi-systemic (have different meanings for stakeholders of different domains) McGuire et al. (2023).
To support advancing systems engineering to higher levels of maturity (Section 5), we now outline key challenges and research opportunities for the modeling community.
@@ 6.1 Process methods
Proper assessment of sustainability requires reasoning about the properties of the engineered system in conjunction with the engineering process (Machado et al., 2020; David et al., 2018). The lack of formal rigor in systems engineering process models (Barisic et al., 2023) renders the end-to-end assessment and optimization of engineering endeavors an insurmountable challenge. While isolated efforts have been made, e.g., in sustainable manufacturing (Andersson et al., 2011) and business information systems (Mihale-Wilson et al., 2022), holistic end-to-end approaches for sustainable systems engineering are lacking. Existing process methods have the potential to be repurposed and tailored for the needs of circular systems engineering. For example, substantial research has been conducted in process methods for finding quality and cost trade-offs in multi-disciplinary design (David et al., 2018). These methods are likely capable to support technical and economic sustainability aspects of systems engineering. Extensions can be made in terms of supported lifecycle phases, e.g., by addressing the post-life of systems, and in terms of externalizing sustainability properties as first-class citizens of complex process models Challenger et al. (2020). For the latter, environmental sustainability should be considered as an immediate target, e.g., building on activity-based cost modeling and value stream mapping coupled with discrete event simulation Andersson et al. (2011). A key challenge to address is the modeling of process networks. Sound composition mechanisms Pratt (1982) could help compose isolated processes into process networks. Current methods for the comprehensive modeling, analysis, and optimization of such process networks are limited. Notably, process monitoring and prediction frameworks that are able to leverage the entirety of enterprise data sources, are completely missing from the state of the art Oberdorf et al. (2023).
@@ 6.2 Modeling and optimizing for sustainability
The bipartite sustainability principle of circular systems engineering (Section 4.2) necessitates proper modeling and simulation methods for assessing the sustainability properties of systems and methods, and to find trade-offs among these properties.
@@
Currently, the evaluation of sustainability in software-intensive systems is mostly a manual effort Venters et al. (2014). There is a pressing need for better automated and more comprehensive methods. As mentioned at the outset, sustainability is a multi-faceted notion involving many different interrelated aspects. This renders the assessment of sustainability particularly challenging. There are only a few domain-specific languages available for specifying sustainability goals Gramelsberger et al. (2023); Lago et al. (2015) despite the ability of modeling to tame complexity. Given the multi-systemic nature of sustainability, multi-paradigm modeling Vangheluwe et al. (2002), multi-view modeling Cicchetti et al. (2019), and user-friendly, flexible modeling approaches, such as blended modeling David et al. (2023b) can provide solid foundations for the next generation of sustainability modeling frameworks and tools. Some of the key challenges in this aspect include the integration of heterogeneous models and the need for personalizing specific views and domain-specific concepts for different stakeholders Combemale et al. (2016). Iterative and incremental evaluation methods Penzenstadler et al. (2021); Gramelsberger et al. (2023) allow for assessing and re-assessing sustainability properties continuously, as systems are developed. However, systematic methods are missing.
Frameworks, such as the Sustainability Awareness Framework (SusAF) Duboc et al. (2019); Brooks et al. (2023) or the Sustainability Evaluation Experience R (SEER) Kienzle et al. (2020), and taxonomies, such as the one by Bischoff et al. (2022) help engineers understand the sustainability impacts of systems, and that, in all relevant sustainability dimensions. However, even such state-of-the-art frameworks fall short of quantifying sustainability and providing actionable insights into sustainability shortcomings. Tasdemir and Gazo (2018) recommend that modeling for sustainability assessment needs to encompass all verticals down to physical processes and that Lean methodologies could set the foundation for comprehensive sustainability assessment frameworks. Such ideas have been explored in, e.g., sustainable supply chain assessment by value stream mapping Sparks and Badurdeen (2014) and in environmental cost modeling through activity based costing Andersson et al. (2011) combined with simulation.
@@
Putting sustainability assessment methods in place enables optimizing systems engineering processes for sustainability goals. Optimization, however, is not trivial given the stratified and multi-systemic nature of sustainability. For example, optimizing towards one SDG target may have negative side-effects on another one Renaud et al. (2022); Tzachor et al. (2022). The general challenge, therefore, is finding acceptable and responsible balancing trade-offs between different sustainability goals. Adopting the mindset of circular systems engineering allows for further opportunities to find trade-offs across different phases of engineering processes. As it is put more explicitly by Tzachor et al. (2022), “in terms of the SDGs, trade-offs are guaranteed to arise given the inherent contradictions across the 169 targets”, and by Zhang et al. (2013)industry is confronted with the challenge of balancing economic and financial priorities against environmental and social responsibilities”. The ability to identify, analyze, and optimize the sustainability trade-offs within and across digital threads by different SDG targets, Key Performance Indicators (KPIs), and Key Environmental Indicators (KEIs) are key to fostering circular systems engineering. Capabilities for the analysis and optimization of single lifecycle phases for often contradictory functional and sustainability goals are required that also translate the rather global and socio-economical metrics like SDG targets and even KPIs toward sustainability-related metrics like KEIs Organisation for Economic Co-operation and Development (2008) (OECD). As a foundation for this trade-off analysis and optimization, methods are required that foster the translation of the generic metrics to fit the context of the current organization and domain.
@@ 6.3 Digital enablers
Digital twins are well-positioned to govern end-to-end processes Heithoff et al. (2023), contributing to important operational goals, such as adaptive control and predictive maintenance Rolnick et al. (2022) for sustainability. However, Tzachor et al. (2022) report multiple challenges that might prevent digital twins to support sustainability goals. The level of digital maturity (or the lack thereof) often hinders proper data management, sometimes to the point where real-time data may not be available or is of poor quality. By McKinsey’s industry digitalization index Harvard Business Review (2016), particularly challenged are agriculture, construction, and healthcare—three sectors strongly linked to sustainable development through precision agriculture, smart cities, and more patient-oriented health systems, respectively. The feasibility of adopting digital twins is also far from being trivial David et al. (2023a), and it is usually a challenge outside of well-developed countries. Furthermore, modeling and simulation of systems with social elements is a wicked problem Coyne (2005), and studies on applying digital twins to social problems are currently lacking. To tackle the challenges of siloed environments and leverage the benefits of digital twinning, interoperability mechanisms of digital twins are to be developed Piroumian (2021). Interoperability allows loosely coupled digital twins to scale up into hierarchies of integrated digital twins.
@@ 7 Conclusion
While the idea of circularity has been around for decades and has been generally recognized as a desirable direction for humankind’s overall sustainability and innovation endeavors, systems engineering has yet to adopt sustainability as a governing principle. With the looming paradigm shift from a technology-focused Industry 4.0 towards a value-focused Industry 5.0, systems engineering now has a rare opportunity to promote sustainability to a first-class citizen. We believe circular systems engineering can be a guiding principle along this road.
To drive the adoption of circularity principles and instigate future research, we have derived prerequisites for circular systems engineering, proposed a sustainability maturity model together with maturity maturation strategies, and identified key challenges and opportunities. Researchers and practitioners can use the paradigm to frame their sustainability-themed research and development activities and to identify the modeling, simulation, and digital transformation goals appropriate to the level of sustainability maturity of their organization.

### 20 Jan 2024 v3

@@ 5.1 Sustainability maturity levels
typically seeks a balance between the quality, cost, and delivery time of engineered systems. At this level of maturity, efforts primarily aim at a better understanding of these three factors in either of the two dimensions of the framework in Figure 2, but not in combination of the two. This typically leads to advanced process methods in the end-to-end dimension of the framework in Figure 2. Substantial work has been dedicated to engineering process modeling Pahl and Beitz (1996), assessment Rout et al. (2007), optimization David et al. (2016), and adaptation Seiger et al. (2019). Occasionally, organizations might realize that quality, cost, and time-to-market can be understood through the lens of sustainability, leading to advancement in the sustainability dimension of the framework. For example, sustainability can be interpreted as a property of systems quality Lago et al. (2015), which allows for a sustainability-driven assessment of quality and cost factors. Ensuring technical sustainability Razavian et al. (2014) also allows for reduced costs by proper evolution after the system has been implemented. Currently, sustainability efforts are mostly focused on energy consumption of software-intensive systems Ournani et al. (2020); Palomba et al. (2019) and advanced digital facilities Bellis and Denil (2022), and on reducing waste of engineering processes Andersson et al. (2011).
@@
combines elements of advanced process methods and extended sustainability scopes. Most importantly, this sustainability maturity level allows reasoning about system-level sustainability along the engineering process and across process networks. This is demonstrated by sustainability-focused end-to-end methods, such as process scheduling for waste management (Le Hesran et al., 2019). Another example is manufacturing rationalization Mouzon et al. (2007); Gao et al. (2018) enabled by an improved understanding of the end-to-end process network and the resulting improved schedulability. Unfortunately, the stratified nature of sustainability McGuire et al. (2023) hinders the development of methods combining process-based longitudinal reasoning with reasoning about sustainability, despite the well-understood benefits (Daoutidis et al., 2016; Barisic et al., 2023). Thus, sustainable systems engineering is an ambitious goal for the majority of organizations currently.
@@
improves on the final limitation of sustainable systems engineering and allows for combining end-to-end concerns with bipartite sustainability. In this combination of concerns, value retention loops are facilitated through the thorough understanding of $(i)$ the network of systems engineering processes, $(ii)$ the sustainability properties of the engineered system, and $(iii)$ the sustainability properties of the employed engineering methods. Circular systems engineering enables assessing and optimizing for systems properties, including sustainability properties across multiple engineering endeavors and allows for finding sustainability trade-offs across different engineered systems. In the most characteristic case, trade-offs can be identified in the upcycling of a systems Wegener (2016)—and that, at the design phase or lifetime of the original system. In heterogeneous cyber-physical systems, this often might be achieved at a relatively low cost by replacing software that operates physical components. For example, older smartphones can be used as cheap sensors Li et al. (2010) to manage the sustainability pressure stemming from the constant need to upgrade computing devices due to Moore’s law. Design for reuse Richardson (2011) and sustainable remanufacturing Jensen et al. (2019) are examples of initiatives that can benefit from a circular systems engineering mindset. Recently, digital twins opened up opportunities for efficient data harvesting and rich decision support based on modeling and simulation Heithoff et al. (2023) and might serve as the enablers of this level of maturity.
@@ 5.2 Sustainability maturation strategies
Maturation strategies from traditional to circular systems engineering can generally follow three pathways.
@@
aims to achieve a better understanding of the value chain and prioritizes improvements in the end-to-end dimension before dedicating efforts to sustainability.
Dahlgaard and Dahlgaard (1999) define operational excellence (OpEx) by the four ‘P’s of excellence: people, partnerships, processes, and products. Covering the end-to-end value chain is of particular importance in achieving operational excellence. Along an operational excellence-driven strategy, a company would develop its internal business capabilities to the level where internal organizations and departments are well-connected through the enterprise process network layer, allowing for the efficient flow of information and advanced analytics techniques such as real-time and predictive analytics Oberdorf et al. (2023) that help achieve excellence.
This strategy might be dominant in sectors where sustainability concerns of engineered systems and, more importantly, engineering methods are dominated by other functional and extra-functional criteria. Such a maturation strategy can be observed in software engineering, where lifecycle models have been researched for decades, but we are just at the beginning of understanding the environmental impacts of computationally demanding methods.
Eventually, an OpEx-driven strategy would implement more elaborate sustainability concerns to arrive at circular systems engineering.
@@
This strategy might be dominant in sectors where sustainability concerns are of particular importance, either due to legislation or social expectations. Pertinent examples include industries with increased environmental impact, such as machine production, cyber-physical systems, or the automotive industry.
@@
aims to balance the previous two strategies, gradually expanding in end-to-end and sustainability concerns.
@@
This strategy might be the most typical and natural as it allows for quick wins at a particular time and context, and the refined scopes of iteration minimize strategic risks.
@@ 6 Challenges and Research Opportunities
System complexity often renders the assessment of sustainability and finding trade-offs among functional and sustainability properties a wicked problem Rittel (1967). Such wicked problems are often impossible to solve due to the complex and interconnected nature of the components of systems and the ensuing sub-problems. For example, of the seventeen SDGs formulated by the UN, fourteen directly incorporate socio-economic elements (goals 1-12, 16, and 17), rendering their modeling and analysis a particularly challenging problem. In such a convoluted problem space, modeling, and especially model-based and model-driven engineering techniques Schmidt (2006), are powerful tools in taming complexity Bork et al. (2023). Through the mechanism of abstraction, model-driven techniques aid understanding of sustainability properties that are often stratified (have different meanings at different levels of abstraction) and multi-systemic (have different meanings for stakeholders of different domains) McGuire et al. (2023).
We now outline key challenges and research opportunities for the modeling community in support of achieving higher sustainability maturity levels (Section 5) in systems engineering.
@@ 6.1 Process methods
Assessing sustainability requires reasoning about the properties of the engineered system in conjunction with the engineering process (Machado et al., 2020; David et al., 2018). The lack of formal rigor in systems engineering process models (Barisic et al., 2023) renders the end-to-end assessment and optimization of engineering endeavors an insurmountable challenge. While isolated efforts have been made, e.g., in sustainable manufacturing (Andersson et al., 2011) and business information systems (Mihale-Wilson et al., 2022), holistic end-to-end approaches for sustainable systems engineering are lacking. Existing process methods have the potential to support the modeling needs of circular systems engineering. For example, substantial research has been conducted on process methods for finding quality and cost trade-offs in multi-disciplinary design (David et al., 2018). These methods might be capable of supporting technical and economic sustainability aspects of systems engineering. Extensions can be made in terms of supported lifecycle phases, e.g., by addressing the post-life of systems and in terms of externalizing sustainability properties as first-class citizens of complex process models Challenger et al. (2020). For the latter, environmental sustainability should be considered as an immediate target, e.g., building on activity-based cost modeling and value stream mapping coupled with discrete event simulation Andersson et al. (2011). A key challenge to address is the modeling of process networks. Sound composition mechanisms Pratt (1982) could help compose isolated processes into process networks. Current methods for the comprehensive modeling, analysis, and optimization of such process networks are limited. Notably, process monitoring and prediction frameworks that are able to leverage the entirety of enterprise data sources are completely missing from the state of the art Oberdorf et al. (2023).
@@ 6.2 Modeling and optimizing for sustainability
The bipartite sustainability principle of circular systems engineering (Section 4.2) necessitates proper modeling and simulation methods for assessing the sustainability properties of systems and methods and finding trade-offs among these properties.
@@
Currently, the evaluation of sustainability in software-intensive systems is mostly a manual effort Venters et al. (2014). There is a pressing need for better automated and more comprehensive methods. As mentioned at the outset, sustainability is a multi-faceted notion involving many different interrelated aspects. This renders the assessment of sustainability challenging. There are only a few domain-specific languages available for specifying sustainability goals Gramelsberger et al. (2023); Lago et al. (2015) despite the ability of modeling to tame complexity. Given the multi-systemic nature of sustainability, multi-paradigm modeling Vangheluwe et al. (2002), multi-view modeling Cicchetti et al. (2019), and user-friendly, flexible modeling approaches, such as blended modeling David et al. (2023b), can provide solid foundations for the next generation of sustainability modeling frameworks and tools. Key challenges in this aspect include the integration of heterogeneous models and the need to personalize specific views and domain-specific concepts for different stakeholders Combemale et al. (2016). Iterative and incremental evaluation methods Penzenstadler et al. (2021); Gramelsberger et al. (2023) allow for assessing and re-assessing sustainability properties continuously as systems are developed. However, systematic methods are missing.
Frameworks, such as the Sustainability Awareness Framework Duboc et al. (2019); Brooks et al. (2023) and taxonomies, such as the one by Bischoff et al. (2022), help engineers understand the sustainability impacts of systems in all relevant sustainability dimensions. However, even state-of-the-art frameworks fall short of quantifying sustainability and providing actionable insights into sustainability shortcomings. Tasdemir and Gazo (2018) recommend modeling for sustainability assessment to encompass all verticals down to physical processes, for which Lean methodologies could set the foundation. Such ideas have been explored in, e.g., sustainable supply chain assessment by value stream mapping Sparks and Badurdeen (2014) and in environmental cost modeling through activity-based costing Andersson et al. (2011).
@@
Putting sustainability assessment methods in place enables optimizing systems engineering processes for sustainability goals. Optimization, however, is not trivial given the stratified and multi-systemic nature of sustainability. For example, optimizing towards one SDG target may have negative side-effects on another one Renaud et al. (2022); Tzachor et al. (2022). The general challenge, therefore, is finding acceptable and responsible balancing trade-offs between different sustainability goals. Adopting the mindset of circular systems engineering allows for further opportunities to find trade-offs across different phases of engineering processes. As put more explicitly by Tzachor et al. (2022), “in terms of the SDGs, trade-offs are guaranteed to arise given the inherent contradictions across the 169 targets”, and by Zhang et al. (2013)industry is confronted with the challenge of balancing economic and financial priorities against environmental and social responsibilities”. The ability to identify, analyze, and optimize the sustainability trade-offs within and across digital threads by different SDG targets, Key Performance Indicators (KPIs), and Key Environmental Indicators (KEIs) are key to fostering circular systems engineering. Capabilities for the analysis and optimization of single lifecycle phases for often contradictory functional and sustainability goals are required that also translate the rather global and socio-economical metrics like SDG targets and even KPIs toward sustainability-related metrics like KEIs Organisation for Economic Co-operation and Development (2008) (OECD). As a foundation for this trade-off analysis and optimization, methods are required that foster the translation of the generic metrics to fit the context of the current organization and domain.
@@ 6.3 Digital enablers
Digital twins are well-positioned to govern end-to-end processes Heithoff et al. (2023), contributing to important operational goals, such as adaptive control and predictive maintenance Rolnick et al. (2022) for sustainability. However, Tzachor et al. (2022) report challenges that might prevent digital twins from supporting sustainability goals. The level of digital maturity (or the lack thereof) often hinders proper data management, sometimes to the point where real-time data may not be available or is of poor quality. According to McKinsey’s industry digitalization index Harvard Business Review (2016), particularly challenged are agriculture, construction, and healthcare—three sectors strongly linked to sustainable development through precision agriculture, smart cities, and more patient-oriented health systems, respectively. The feasibility of adopting digital twins is also far from being trivial David et al. (2023a), and it is usually a challenge outside of well-developed countries. Furthermore, modeling and simulation of systems with social elements is a wicked problem Coyne (2005), and studies on applying digital twins to social problems are currently lacking. To tackle the challenges of siloed environments and leverage the benefits of digital twinning, interoperability mechanisms of digital twins are to be developed Piroumian (2021) to allow for loosely coupled digital twins to scale up into robust hierarchies.
@@ 7 Conclusion
While the idea of circularity has been around for decades and is recognized as a desirable direction for humankind’s overall sustainability and innovation endeavors, systems engineering has yet to adopt sustainability as a governing principle. With the looming paradigm shift from a technology-focused Industry 4.0 towards a value-focused Industry 5.0, systems engineering now has a rare opportunity to promote sustainability to a first-class citizen. We believe circular systems engineering can be a guiding principle along this road.
To drive the adoption of circularity principles and instigate future research, we have derived prerequisites for circular systems engineering, proposed a sustainability maturity model together with maturity maturation strategies, and identified key challenges and opportunities. Researchers can use the paradigm to contextualize their sustainability-themed research. Practitioners can rely on the presented maturity framework and maturation strategies to identify the modeling, simulation, and digital transformation goals appropriate to their organizations’ sustainability ambitions.